A college student earned $8100 during the summer. They invested part of the money at 8% and the rest at 6%. If they received a total of $564 in interest at the end of the year how much was invested at 8%? 
Answer by mananth(16946)

Investment I 8.00% per annum ----x
Investment II 6.00% per annum ----y
The sum of the investments is $8,100.00
The sum of individual interests = $564.00
x + y 8100 ------------------------1
8.00% x + 6.00% y = $564.00
Multiply by 100
8 x + 6 y = $56,400.00--------2

Multiply (1) by -8
we get

-8x-8y= -64800.00

Add this to (2)

0x-2y=-$8,400.00

divide by-2

y = $4,200.00 investment at 6.00%
Balance $3,900.00 investment at 8.00%
CHECK
$3,900.00 $312.00
$4,200.00 $252.00
Total $564.00